### v1

# Optimal Best-Arm Identification under Fixed Confidence with Multiple Optima

Abstract
We study the problem of best-arm identification in stochastic multi-armed bandits under the fixed-confidence setting, with a particular focus on instances that admit multiple optimal arms. While the Track-and-Stop algorithm of Garivier and Kaufmann (2016) is widely conjectured to be instance-optimal, its performance in the presence of multiple optima has remained insufficiently understood. In this work, we revisit the Track-and-Stop strategy and propose a modified stopping rule that ensures instance-optimality even when the set of optimal arms is not a singleton. Our analysis introduces a new information-theoretic lower bound that explicitly accounts for multiple optimal arms, and we demonstrate that our stopping rule tightly matches this bound.
Index Terms:
Multi-Armed Bandits, Best-Arm Identification, Sample Complexity.
I Introduction
The multi-armed bandit (MAB) problem is a classic framework for sequential decision making under uncertainty, where a learner interacts with a set of arms representing unknown reward distributions. The objective is to gather information about these arms to make an informed decision.
In the best-arm identification (BAI) problem under the fixed-confidence setting, the learner’s goal is to identify an arm whose mean is maximal (or nearly so) with a prescribed confidence level, using as few samples as possible. This setting has broad applications in areas such as clinical trials, A/B testing, hyperparameter tuning, and recommendation systems.
While much of the prior work on BAI focuses on the case where there is a unique best arm, many practical scenarios involve multiple optimal arms—that is, more than one arm may achieve the highest expected reward. This work addresses a fundamental gap in the literature by developing a theoretical and algorithmic framework that achieves optimal sample complexity in the presence of multiple optimal arms.
I-A Motivation
In best-arm identification problems, the primary goal is to identify an optimal arm from a set of stochastic options using as few samples as possible while maintaining a fixed confidence level. This problem is fundamental to sequential decision-making and has broad applicability in domains such as clinical trials, A/B testing, and hyperparameter tuning. While much of the existing work focuses on settings with a unique best arm, many real-world applications naturally involve multiple arms achieving the same maximal expected reward. This motivates the need to study best-arm identification in the more general scenario where the set of optimal arms is not a singleton.
Among algorithms proposed for fixed-confidence best-arm identification, the Track-and-Stop algorithm stands out due to its principled design based on likelihood ratio statistics and its strong empirical performance. Although it is widely believed to be near-optimal, especially in the case of a single best arm, a complete theoretical justification of its optimality—particularly in instances with multiple optimal arms—has remained an open question.
In this work, we address this gap by modifying the stopping rule of the Track-and-Stop algorithm and proving that the resulting algorithm achieves instance-optimal sample complexity in the fixed-confidence setting, even in the presence of multiple optimal arms. Our analysis introduces a refined information-theoretic lower bound tailored to this general setting and shows that the proposed stopping condition tightly matches this bound. This result provides the first formal guarantee of optimality for Track-and-Stop with multiple optima and offers new insights into the design of exploration strategies in stochastic bandit problems.
I-B Related Work
The problem of best-arm identification (BAI) in stochastic multi-armed bandits has been a central focus in the bandit literature for over a decade, particularly under the fixed-confidence and fixed-budget settings. In the fixed-confidence framework, the goal is to identify an arm with the highest expected reward with probability at least $1-\delta$, while minimizing the expected number of samples (even2006action, ; garivier2016optimal, ). The fixed-budget setting, on the other hand, fixes the total sampling budget and aims to minimize the error probability (gabillon2012best, ).
Early foundational work in fixed-confidence BAI introduced elimination strategies and stopping criteria that guarantee PAC (Probably Approximately Correct) bounds (even2006action, ). These were later complemented by information-theoretic approaches that established matching lower and upper bounds on sample complexity under the assumption of a unique best arm (garivier2016optimal, ; garivier2016track, ). Such analyses rely on likelihood ratio statistics and characterize the complexity of the problem through Kullback-Leibler (KL) divergences between the reward distributions of arms. Algorithms such as Track-and-Stop (garivier2016optimal, ) and lil’UCB (jamieson2014lil, ) leverage these theoretical insights to achieve near-optimal sample efficiency.
Algorithms like LUCB (kalyanakrishnan2012pac, ) and lil’UCB adapt confidence bounds to balance exploration across arms efficiently, achieving PAC guarantees with sample complexities that depend on the inverse squared gaps between arm means. Despite their success, these methods generally rely on the uniqueness of the best arm, and their performance can degrade when multiple arms share the highest mean reward.
In many practical applications—such as clinical trials, recommendation systems, and A/B testing—multiple arms may be equally optimal due to noise or inherent problem structure. The presence of multiple optimal arms poses unique challenges because standard algorithms continue to sample these tied arms excessively to differentiate among them, leading to unnecessary increases in sample complexity.
The literature addressing multiple optimal arms has mostly focused on related but distinct problems such as top-$k$ arm identification or best subset selection (gabillon2012best, ; chen2014combinatorial, ; jamieson2014bandit, ). These works aim to identify the entire set of best arms with high confidence, often requiring more samples than necessary if the goal is only to find any one best arm. Some recent studies have begun to recognize this inefficiency; for instance, (qin2022gap, ) propose gap-free algorithms that adapt automatically to the presence of zero gaps (i.e., ties) and reduce redundant exploration. Similarly, (combes2021minimal, ) analyze minimal exploration strategies and highlight how ignoring ties can cause suboptimal sampling strategies.
More recent advances by (hu2023almost, ) propose algorithms explicitly designed to exploit the existence of multiple optimal arms. These algorithms feature stopping rules and sampling strategies that avoid unnecessary comparisons between tied arms and achieve nearly instance-optimal sample complexity under fixed confidence. Their approach shows significant improvements over classical methods when multiple arms are truly optimal.
From a theoretical standpoint, classical lower bounds on sample complexity assume uniqueness of the best arm (garivier2016optimal, ). The multiple-optimal-arm scenario requires refined lower bounds that reflect the complexity of identifying any one of the optimal arms, demonstrating how sampling effort should be distributed among tied arms rather than wasted on differentiating them. This nuanced understanding informs the design of more efficient algorithms tailored to the presence of ties.
Beyond fixed-confidence BAI, related problems such as fixed-budget best-arm identification also face challenges with multiple optimal arms. Studies like (gabillon2012best, ) and (chen2014combinatorial, ) investigate how limited sampling budgets affect the ability to identify top arms accurately, with recent insights showing that tie-aware sampling can improve error rates. Furthermore, problems involving combinatorial bandits (chen2014combinatorial, ), contextual bandits (simchi2021best, ), and reinforcement learning where multiple optimal policies exist also benefit from algorithms capable of recognizing and exploiting ties among arms or actions.
Besides efficiency considerations, the presence of multiple optimal arms has practical implications for fairness and diversity. When several arms represent different treatments, products, or policies with equivalent expected performance, selecting any one of them ensures equitable treatment and avoids unintended bias in recommendations or decisions (xu2021fairness, ). Algorithms that identify ties quickly can facilitate fairer deployment strategies in real-world applications.
In summary, while much of the classical best-arm identification literature focuses on unique optima, recent research increasingly recognizes the importance of handling multiple optimal arms. Our work contributes to this growing area by refining the Track-and-Stop algorithm through a new stopping condition that is both tie-aware and theoretically grounded. We prove that this modified version achieves instance-optimal sample complexity in the fixed-confidence setting, even when the set of optimal arms is not a singleton. Our approach closes a fundamental theoretical gap in the literature and offers new insights into designing efficient, fair, and practically relevant algorithms for modern stochastic bandit problems.
I-C Contributions
This work makes the following key contributions:
  • Modified Track-and-Stop with Optimal Stopping: We propose a refined version of the Track-and-Stop algorithm featuring a novel stopping rule that achieves instance-optimal sample complexity for best-arm identification under the fixed-confidence setting, even when multiple optimal arms exist. This resolves a long-standing open problem in the literature.
  • New Information-Theoretic Lower Bound: We introduce a novel lower bound on sample complexity that explicitly accounts for multiple optimal arms, characterizing the minimal number of samples required to identify any optimal arm with high confidence—avoiding unnecessary distinctions between tied arms.
  • Refined Theoretical Analysis: We develop an analytical framework that combines generalized likelihood ratio statistics with convex optimization techniques, enabling a tight alignment between the proposed stopping criterion and the information-theoretic lower bounds.
  • Adaptivity to Multiple Optimal Arms: Our analysis covers settings with multiple optimal arms, showing that the modified Track-and-Stop algorithm automatically adapts to such structures and preserves optimal performance guarantees.
I-D Notations
For two distributions $P$ and $Q$ defined over a common support $\mathcal{X}$, the Kullback–Leibler (KL) divergence from $P$ to $Q$ is given by
[…]
where $\frac{dP}{dQ}$ denotes the Radon–Nikodym derivative of $P$ with respect to $Q$. This formulation is valid for both continuous and discrete probability measures.
For any positive real number $x\in\mathbb{R}_{+}$, we use $\log(x)$ to denote the logarithm base 2. Furthermore, define the binary KL divergence function between two Bernoulli parameters $x,y\in[0,1]$ as
$\mathrm{kl}(x,y)=x\log\frac{x}{y}+(1-x)\log\frac{1-x}{1-y},$
with the conventions $\mathrm{kl}(0,0)=\mathrm{kl}(1,1)=0$.
For any functions $f:\mathbb{R}\to\mathbb{R}$, we define
$\dot{f}(x)=\frac{df(x)}{dx},\qquad\ddot{f}(x)=\frac{d^{2}f(x)}{dx^{2}}.$
Finally, let
$\Sigma_{K}=\left\{w\in\mathbb{R}_{+}^{K}:\sum_{i=1}^{K}w_{i}=1\right\}$
denote the $K$-dimensional probability simplex.
II One-parameter exponential family
The exponential family encompasses a broad and significant class of probability distributions, including many widely used models in statistics and machine learning, such as the Bernoulli, Poisson, Gaussian (with known variance), and exponential distributions. A particularly important subclass is the one-parameter exponential family, where each distribution is characterized by a single real-valued parameter.
A probability distribution $P_{\mu}$ belongs to the one-parameter exponential family if its density (or mass) function can be written in the form:
$p_{\mu}(x)=\exp\left(\theta(\mu)x-A(\theta(\mu))\right),$ (1)
where:
  • $\mu=\mathbb{E}[X]$ denotes the mean of the distribution,
  • $\theta(\mu)$ is the natural (canonical) parameter, assumed to be a bijective function of $\mu$,
  • $A(\theta)$ is the log-partition function, which is convex and differentiable, and ensures normalization:
[…]
  • $d(\mu,\lambda):=D_{\mathrm{KL}}(P_{\mu}\|P_{\lambda})$ denotes the Kullback–Leibler (KL) divergence between two members of the family.
The KL divergence between $P_{\mu}$ and $P_{\lambda}$ admits the following closed-form expression:
$d(\mu,\lambda)=A(\theta(\lambda))-A(\theta(\mu))-(\theta(\lambda)-\theta(\mu))\mu.$ (2)
Then, the following results can be proved.
Lemma 1.
For the one-parameter exponential family, it holds that
$\frac{d\theta(\mu)}{d\mu}=\frac{1}{\ddot{A}(\theta(\mu))}>0,\qquad\forall\mu.$ (3)
Proof.
The mean of the distribution is given by:
$\mu(\theta)=\mathbb{E}_{\theta}[X]=\dot{A}(\theta).$ (4)
Since the function $A(\theta)$ is strictly convex, its derivative $\dot{A}(\theta)$, which maps $\theta\mapsto\mu$, is strictly increasing.
From calculus, if a function $f$ is strictly increasing and differentiable, then its inverse $f^{-1}$ is differentiable, and
$\frac{d}{dy}f^{-1}(y)=\frac{1}{\dot{f}(f^{-1}(y)}.$ (5)
Applying this to $f(\theta)=\mu(\theta)=\dot{A}(\theta)$, whose inverse is $\theta(\mu)$, we get:
$\frac{d}{d\mu}\theta(\mu)=\frac{1}{\ddot{A}(\theta(\mu))}.$ (6)
Because $A(\theta)$ is the log-partition function of an exponential family, it is strictly convex, and hence:
$\ddot{A}(\theta)>0,\qquad\forall\theta.$ (7)
Therefore:
$\frac{d}{d\mu}\theta(\mu)=\frac{1}{\ddot{A}(\theta(\mu))}>0,\qquad\forall\mu.$ (8)
The following result can be shown in Appendix A.
Lemma 2.
Let $(w_{1},w_{2},\cdots,w_{M},w_{a})$ be a tuple of positive numbers. For common exponential family distributions such as Gaussian with known variance, Bernoulli, and Poisson, it holds that
$f(\lambda_{1},\lambda_{a})=\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a})$
is jointly convex in $(\lambda_{1},\lambda_{2},\cdots,\lambda_{M},\lambda_{a})$, where $d(\mu,\lambda)$ is the KL divergence between two distributions from the exponential family defined in (2).
III Problem Setup
We consider the stochastic multi-armed bandit (MAB) problem in the fixed-confidence best-arm identification setting. Let $\mathcal{A}=\{1,2,\dots,K\}$ be a finite set of $K$ arms. Each arm $a\in\mathcal{A}$ is associated with an unknown distribution $\nu_{a}$ belonging to a known exponential family, with mean reward $\mu_{a}=\mathbb{E}_{X\sim\nu_{a}}[X]$.
At each round $t\in\mathbb{N}$, the learner selects an arm $A_{t}\in\mathcal{A}$ and observes an independent sample $X_{t}\sim\nu_{A_{t}}$. Based on the history of arm pulls and rewards, i.e., $\mathcal{F}_{t}=\sigma(A_{1},X_{1},A_{2},X_{2},\cdots,A_{t},X_{t})$, the learner chooses a stopping time $\tau$ and outputs a recommendation $\hat{a}_{\tau}\in\mathcal{A}$.
Define the set of optimal arms as
$\mathcal{A}^{\star}:=\left\{a\in\mathcal{A}\,:\,\mu_{a}=\mu^{\star}\right\},\quad\text{where}\quad\mu^{\star}:=\max_{a\in\mathcal{A}}\mu_{a}.$
The goal is to identify any one arm in $\mathcal{A}^{\star}$ with high probability while minimizing the expected sample complexity.
Fixed-Confidence Guarantee.
Given a confidence parameter $\delta\in(0,1)$, a $(\delta,\mathcal{A}^{\star})$-PAC (Probably Approximately Correct) algorithm is one that ensures
$\mathbb{P}(\hat{a}_{\tau}\in\mathcal{A}^{\star})\geq 1-\delta.$
Sample Complexity.
The performance of an algorithm is measured by its sample complexity, i.e., the expected number of samples used before stopping: $\mathbb{E}[\tau].$ Our aim is to design algorithms that satisfy the PAC criterion and achieve the minimal possible sample complexity, especially in the presence of multiple optimal arms.
Challenge with Multiple Optima.
When $|\mathcal{A}^{\star}|>1$, standard algorithms may over-sample to distinguish between arms that are equally optimal. The key challenge is to avoid unnecessary comparisons among optimal arms and stop as soon as any one of them is confidently identified.
IV General Lower Bound
First, we recall the following lemma.
Lemma 3.
(Kaufmann2014OnTC, , Lemma 1) Let $\nu$ and $\nu^{\prime}$ be two bandit models with $K$ arms such that for all $a$, the distributions $\nu_{a}$ and $\nu^{\prime}_{a}$ are mutually absolutely continuous. For any almost-surely finite stopping time $\sigma$ with respect to $(\mathcal{F}_{t})$,
$\sum_{a=1}^{K}\mathbb{E}_{\nu}[N_{a}(\sigma)]D_{\text{KL}}(\nu_{a},\nu^{\prime}_{a})\geq\sup_{\mathcal{E}\in\mathcal{F}_{\sigma}}\text{kl}(\mathbb{P}_{\nu}(\mathcal{E}),\mathbb{P}_{\nu^{\prime}}(\mathcal{E})).$ (9)
Denote
$\mathcal{S}=\big{\{}\mu:\mu_{1}=\mu_{2}=\cdots=\mu_{M}>\mu_{M+1}\geq\mu_{M+2}\geq\cdots\geq\mu_{K}\big{\}}.$ (10)
Then, the following can be proved.
Theorem 4.
For any tuple $(\alpha_{1},\alpha_{2},\cdots,\alpha_{M})\in[0,1]^{M}$ such that $0\leq\sum_{i=1}^{M}\alpha_{i}\leq 1$, define
$I_{\alpha_{1},\alpha_{2},\cdots,\alpha_{M}}(\mu_{1},\mu_{2},\cdots,\mu_{M},\mu_{a})$ $=\sum_{i=1}^{M}\alpha_{i}d\bigg{(}\mu_{i},\sum_{i=1}^{M}\alpha_{i}\mu_{i}+\bigg{(}1-\sum_{i=1}^{M}\alpha_{i}\bigg{)}\mu_{a}\bigg{)}$
$\qquad+\bigg{(}1-\sum_{i=1}^{M}\alpha_{i}\bigg{)}d\bigg{(}\mu_{a},\sum_{i=1}^{M}\alpha_{i}\mu_{i}+\bigg{(}1-\sum_{i=1}^{M}\alpha_{i}\bigg{)}\mu_{a}\bigg{)}.$ (11)
Let $\delta\in(0,1)$. For any $\delta$-PAC strategy and any bandit model $\mu\in\mathcal{S}$,
$\mathbb{E}_{\mu}[\tau_{\delta}]\geq T^{*}(\mu)\text{kl}(1-\delta\,\|\,\delta),$ (12)
where
$T^{*}(\mu)^{-1}=\sup_{w\in\Sigma_{K}}\min_{a\notin\{1,2,\cdots,M\}}\bigg{(}\sum_{i=1}^{M}w_{i}+w_{a}\bigg{)}$
×$\qquad\times I_{\frac{w_{1}}{\sum_{i=1}^{M}w_{i}+w_{a}},\frac{w_{2}}{\sum_{i=1}^{M}w_{i}+w_{a}},\cdots,\frac{w_{M}}{\sum_{i=1}^{M}w_{i}+w_{a}}}(\mu_{1},\mu_{2},\cdots,\mu_{M},\mu_{a}).$ (13)
Remark 5.
As $\text{kl}(1-\delta\,\|\,\delta)\sim\log(1/\delta)$ when $\delta$ goes to zero. Theorem 4 provides an asymptotic lower bound
$\liminf_{\delta\to 0}\frac{\mathbb{E}_{\mu}[\tau_{\delta}]}{\log(1/\delta)}\geq T^{*}(\mu).$ (14)
A non-asymptotic version can be obtained by using the inequality $\text{kl}(1-\delta\,\|\,\delta)\geq\log(1/(2.4\delta))$ that holds for any $\delta\in(0,1)$ (Garivier2016OptimalBA, ).
Proof.
Denote by
$\text{Alt}(\mu)=\bigcup_{a\in\mathcal{A}\setminus\{1,2,\cdots,M\}}\big{\{}\lambda\in\mathcal{S}:(\lambda_{a}>\lambda_{1})\wedge(\lambda_{a}>\lambda_{2})\wedge\cdots\wedge(\lambda_{a}>\lambda_{M})\}.$ (15)
For any $\lambda\in\text{Alt}(\mu)$, there exists $a\notin\{1,2,\cdots,M\}$ such that $\lambda_{a}>\lambda_{i}$ for all $i\in[M]$. Hence, $\{1,2,\cdots,M\}$ is not the set of optimal arms under $\lambda$. Thus, introducing the event =$\mathcal{E}=\{\hat{a}_{\tau}\in\{1,2,\cdots,M\}\}\in\mathcal{F}_{\tau}$, any $\delta$-PAC algorithm satisfies $\mathbb{P}_{\mu}(\mathcal{E})>1-\delta$ and $\mathbb{P}_{\lambda}(\mathcal{E})\leq\delta$.
By Lemma 3, we have
$\forall\lambda\in\text{Alt}(\mu):\sum_{a=1}^{K}d(\mu_{a},\lambda_{a})\mathbb{E}_{\mu}[N_{a}(\tau_{\delta})]\geq\text{kl}(1-\delta,\delta).$ (16)
Then, it holds that
$\text{kl}(\delta,1-\delta)\leq\mathbb{E}_{\mu}[\tau_{\delta}]\sup_{w\in\Sigma_{K}}\inf_{\lambda\in\text{Alt}(\mu)}\bigg{(}\sum_{a=1}^{K}w_{a}d(\mu_{a},\lambda_{a})\bigg{)},$ (17)
or
$\mathbb{E}_{\mu}[\tau_{\delta}]\geq\bigg{[}\sup_{w\in\Sigma_{K}}\inf_{\lambda\in\text{Alt}(\mu)}\bigg{(}\sum_{a=1}^{K}w_{a}d(\mu_{a},\lambda_{a})\bigg{)}\bigg{]}^{-1}\text{kl}(1-\delta,\delta).$ (18)
Now, we estimate $\inf_{\lambda\in\text{Alt}(\mu)}\big{(}\sum_{a=1}^{K}w_{a}d(\mu_{a},\lambda_{a})\big{)}$ for a fixed $w\in\Sigma_{K}$. Observe that
$\inf_{\lambda\in\text{Alt}(\mu)}\bigg{(}\sum_{a=1}^{K}w_{a}d(\mu_{a},\lambda_{a})\bigg{)}$
$\qquad=\min_{a\notin\{1,2,\cdots,M\}}\inf_{\lambda:\lambda_{a}\geq\lambda_{1},\lambda_{a}\geq\lambda_{2},\cdots,\lambda_{a}\geq\lambda_{M}}\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a}).$ (19)
Next, we minimize $f(\lambda_{1},\lambda_{2},\cdots,\lambda_{M},\lambda_{a})=\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a})$ under constraints $\lambda_{a}\geq\lambda_{1},\lambda_{a}\geq\lambda_{2},\cdots,\lambda_{a}\geq\lambda_{M}$. This is a convex optimization problem (see Lemma 2) that can solved analytically.
The Lagrangian is
$L(\lambda_{1},\lambda_{2},\cdots,\lambda_{M},\lambda_{a},\beta_{1},\beta_{2},\cdots,\beta_{M})=\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a})-\sum_{i=1}^{M}\beta_{i}(\lambda_{a}-\lambda_{i}),$ (20)
where $\beta_{i}\geq 0$ for all $i\in[M]$.
For a one-parameter exponential family with density
$p_{\mu}(x)=h(x)\exp\left(\theta(\mu)x-A(\theta(\mu))\right),$
where $\mu=A^{\prime}(\theta)$, the KL divergence between $P_{\mu}$ and $P_{\lambda}$ is given by:
$d(\mu,\lambda)=A(\theta(\lambda))-A(\theta(\mu))-(\theta(\lambda)-\theta(\mu))\mu.$
Hence, we have
$L(\lambda_{1},\lambda_{2},\cdots,\lambda_{M},\lambda_{a},\beta_{1},\beta_{2},\cdots,\beta_{M})$
$\qquad=\sum_{i=1}^{M}w_{i}\big{[}A(\theta(\lambda_{i}))-A(\theta(\mu_{i}))-(\theta(\lambda_{i})-\theta(\mu_{i}))\mu_{i}\big{]}$
$\qquad\qquad+w_{a}\big{[}A(\theta(\lambda_{a}))-A(\theta(\mu_{a}))-(\theta(\lambda_{a})-\theta(\mu_{a}))\mu_{a}\big{]}-\sum_{i=1}^{M}\beta_{i}(\lambda_{a}-\lambda_{i}).$ (21)
Then, the KKT conditions are
00 0 $=\frac{\partial L}{\partial\lambda_{i}}=w_{i}(\lambda_{i}-\mu_{i})\frac{d\theta(\lambda_{i})}{d\lambda_{i}}+\beta_{i},\qquad\forall i\in[M],$ (22)
00 0 $=\frac{\partial L}{\partial\lambda_{a}}=w_{a}(\lambda_{a}-\mu_{a})\frac{d\theta(\lambda_{a})}{d\lambda_{a}}-\sum_{i=1}^{M}\beta_{i},$ (23)
$\beta_{i}(\lambda_{a}-\lambda_{i})$ $=0,\qquad\forall i\in[M],$ (24)
$\lambda_{a}$ $\geq\lambda_{i},\qquad\forall i\in[M].$ (25)
Now, we consider cases:
  • Case 1: $\beta_{i}=0$ for all $i\in[M]$. Then, from (22), (23), and Lemma 1 we obtain $\lambda_{i}=\mu_{i}\quad\forall i\in[M]$ and $\lambda_{a}=\mu_{a}$. Then, we have $\lambda_{a}<\lambda_{i}$ for all $i\in[M]$, which contradicts (25).
  • Case 2: There exists exactly $L$ indices $\{i_{1},i_{2},\cdots,i_{L}\}$ for $1\leq L<M$ such that $\lambda_{a}=\lambda_{i_{k}}$ for all $k\in[L]$. Then, we have
    $w_{i_{k}}(\lambda_{i_{k}}-\mu_{i_{k}})\frac{d\theta(\lambda_{i_{k}})}{d\lambda_{i_{k}}}+\beta_{i_{k}}$ $=0,\qquad\forall k\in[L],$ (26)
    $w_{a}(\lambda_{a}-\mu_{a})\frac{d\theta(\lambda_{a})}{d\lambda_{a}}-\sum_{k=1}^{L}\beta_{i_{k}}$ $=0$ (27)
    It follows that
    $\sum_{k=1}^{L}w_{i_{k}}(\lambda_{i_{k}}-\mu_{i_{k}})\frac{d\theta(\lambda_{i_{k}})}{d\lambda_{i_{k}}}+w_{a}(\lambda_{a}-\mu_{a})\frac{d\theta(\lambda_{a})}{d\lambda_{a}}=0,$ (28)
    which leads to (by combining with Lemma 1)
    $\lambda_{a}=\lambda_{i_{1}}=\cdots=\lambda_{i_{L}}=\frac{\sum_{k=1}^{L}w_{i_{k}}\mu_{i_{k}}+w_{a}\mu_{a}}{\sum_{k=1}^{L}w_{i_{k}}+w_{a}}.$ (29)
    Now, since $L<M$, from (24) there exists $j\in[M]\setminus\{i_{1},i_{2},\cdots,i_{L}\}$ such that $\beta_{j}=0$. Hence, from (22) we have
    $\lambda_{j}=\mu_{j}.$ (30)
    Now, from (29) and (30) we have $\lambda_{a}<\lambda_{j},$ which contradicts (25).
  • Case 3: $\lambda_{1}=\lambda_{2}=\cdots=\lambda_{M}=\lambda_{a}$. Then, from (22) and (23) and using the same arguments as Case 2, we have
    $\lambda_{1}=\lambda_{2}=\cdots=\lambda_{M}=\lambda_{a}=\frac{\sum_{i=1}^{M}w_{i}\mu_{i}+w_{a}\mu_{a}}{\sum_{i=1}^{M}w_{i}+w_{a}}.$ (31)
    Then, we have
    $\min_{a\notin\{1,2,\cdots,M\}}\inf_{\lambda:\lambda_{a}\geq\lambda_{1},\lambda_{a}\geq\lambda_{2},\cdots,\lambda_{a}\geq\lambda_{M}}\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a})$
    $=\min_{a\notin\{1,2,\cdots,M\}}\bigg{(}\sum_{i=1}^{M}w_{i}+w_{a}\bigg{)}I_{\frac{w_{1}}{\sum_{i=1}^{M}w_{i}+w_{a}},\frac{w_{2}}{\sum_{i=1}^{M}w_{i}+w_{a}},\cdots,\frac{w_{M}}{\sum_{i=1}^{M}w_{i}+w_{a}}}(\mu_{1},\mu_{2},\cdots,\mu_{M},\mu_{a}).$ (32)
Finally, by letting
$T^{*}(\mu)^{-1}$ $=\sup_{w\in\Sigma_{K}}\min_{a\notin\{1,2,\cdots,M\}}\inf_{\lambda:\lambda_{a}\geq\lambda_{1},\lambda_{a}\geq\lambda_{2},\cdots,\lambda_{a}\geq\lambda_{M}}\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a}),$ (33)
and
$w^{*}(\mu)=\text{argmax}_{w\in\Sigma_{K}}\min_{a\notin\{1,2,\cdots,M\}}\inf_{\lambda:\lambda_{a}\geq\lambda_{1},\lambda_{a}\geq\lambda_{2},\cdots,\lambda_{a}\geq\lambda_{M}}\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a}),$ (34)
we conclude our proof of this lemma. ∎
V The Track-and-Stop Strategy
V-A Sampling Rule: Tracking the Optimal Proportions
We adopt the same sampling rules as described in (Garivier2016OptimalBA, , Section 3). The core idea is to match the optimal sampling proportions $w^{*}(\mu)$ by tracking the plug-in estimate $w^{*}(\hat{\mu}(t))$. However, in bandit settings, relying directly on plug-in estimates can be risky: poor initial estimates may cause some arms to be prematurely discarded, thus preventing the collection of new observations that would correct earlier mistakes. In fact, naive plug-in tracking may fail to identify the best arm reliably.
To mitigate this issue, a simple yet effective solution is to enforce sufficient exploration across all arms to ensure fast convergence of the empirical means $\hat{\mu}(t)$.
One such strategy, termed C-Tracking, modifies the optimization step by projecting $w^{*}(\mu)$ onto a truncated simplex to ensure every arm is sampled at least minimally. For each $\varepsilon\in(0,1/K]$, define the truncated simplex:
$\Sigma_{K}^{\varepsilon}=\left\{(w_{1},\dots,w_{K})\in[\varepsilon,1]^{K}:\sum_{a=1}^{K}w_{a}=1\right\}.$ (35)
Let $w^{\varepsilon}(\mu)$ denote the $L^{\infty}$-projection of $w^{*}(\mu)$ onto $\Sigma_{K}^{\varepsilon}$, and choose $\varepsilon_{t}=\frac{1}{2\sqrt{K^{2}+t}}$. Then the sampling rule is given by
$A_{t+1}\in\arg\max_{1\leq a\leq K}\sum_{s=0}^{t}w_{a}^{\varepsilon_{s}}(\hat{\mu}(s))-N_{a}(t).$ (36)
An alternative, often more practical rule is D-Tracking, which targets $w^{*}(\hat{\mu}(t))$ directly and introduces forced exploration only when necessary. Let
$U_{t}=\left\{a\in\{1,\dots,K\}:N_{a}(t)<\sqrt{t}-\frac{K}{2}\right\}.$ (37)
Then the sampling rule is defined as
$A_{t+1}\in\begin{cases}\arg\min_{a\in U_{t}}N_{a}(t)&\text{if }U_{t}\neq\emptyset\quad\text{(\emph{forced exploration})},\\ \arg\max_{1\leq a\leq K}\left[tw^{*}_{a}(\hat{\mu}(t))-N_{a}(t)\right]&\text{otherwise (\emph{direct tracking})}.\end{cases}$ (38)
Both sampling rules are guaranteed to ensure that the empirical sampling proportions converge to the optimal allocation, as established in (Garivier2016OptimalBA, , Section 3.1).
V-B Chernoff’s Stopping Rule
Let
$Z_{a;b_{1},b_{2},\cdots,b_{M}}(t)=\log\frac{\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{2}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}{\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})},$ (39)
where
$p_{\mu}(Z_{1},Z_{2},\cdots,Z_{n})=\prod_{k=1}^{n}\exp\big{(}\dot{A}^{-1}(\mu)Z_{k}-A(\dot{A}^{-1}(\mu))\big{)}.$ (40)
Observe that
$p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})$ $=\prod_{k=1}^{N_{a}(t)}\exp\bigg{(}\dot{A}^{-1}(\mu^{\prime}_{a})X_{k}-A(\dot{A}^{-1}(\mu^{\prime}_{a}))\bigg{)}$ (41)
$=\exp\bigg{[}N_{a}(t)\bigg{(}\dot{A}^{-1}(\mu^{\prime}_{a})\hat{\mu}_{a}(t)-A(\dot{A}^{-1}(\mu^{\prime}_{a}))\bigg{)}\bigg{]}.$ (42)
Here,
$\hat{\mu}_{a}(t)=\frac{1}{N_{a}(t)}\sum_{k=1}^{N_{a}(t)}X_{k}.$ (43)
Similarly, we have
$p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})$ $=\exp\bigg{[}N_{b_{i}}(t)\bigg{(}\dot{A}^{-1}(\mu^{\prime}_{b_{i}})\hat{\mu}_{b_{i}}(t)-A(\dot{A}^{-1}(\mu^{\prime}_{b_{i}}))\bigg{)}\bigg{]},\qquad\forall i\in[M].$ (44)
Hence, we have
$\log\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})$
$\qquad=\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}\bigg{[}l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg{]},$ (45)
where
$l_{x}(\mu^{\prime}_{x})=N_{x}(t)\bigg{(}\dot{A}^{-1}(\mu^{\prime}_{x})\hat{\mu}_{x}(t)-A(\dot{A}^{-1}(\mu^{\prime}_{x}))\bigg{)},\qquad\forall x\in\{a,b_{1},b_{2},\cdots,b_{M}\},$ (46)
is concave in $\mu^{\prime}_{x}$.
Now, we have
$\frac{\partial l_{x}(\mu^{\prime}_{x})}{\partial\mu^{\prime}_{x}}$ $=\frac{\partial l_{x}(t)}{\partial\dot{A}^{-1}(\mu^{\prime}_{x})}\frac{\partial\dot{A}^{-1}(\mu^{\prime}_{x})}{\partial\mu^{\prime}_{x}}$ (47)
$=N_{x}(t)(\hat{\mu}_{x}(t)-\mu^{\prime}_{x})\frac{\partial\dot{A}^{-1}(\mu^{\prime}_{x})}{\partial\mu^{\prime}_{x}}$ (48)
$=-N_{x}(t)(\hat{\mu}_{x}(t)-\mu^{\prime}_{x})\bigg{(}\frac{\ddot{A}(\mu^{\prime}_{x})}{(\dot{A}(\mu^{\prime}_{x}))^{2}}\bigg{)},$ (49)
which is equal to zero if and only if $\mu^{\prime}_{x}=\hat{\mu}_{x}(t)$.
This means that
$\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}\bigg{[}l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg{]}\geq l_{a}(\hat{\mu}_{a}(t))+\sum_{i=1}^{M}l_{b_{i}}(\hat{\mu}_{b_{i}}(t)).$ (50)
Under the condition that $\hat{\mu}_{a}(t)\leq\hat{\mu}_{b_{i}}(t)),\forall i\in[M]$, the equality in (50) is achieved, which means that
$\log\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})$
$\qquad=l_{a}(\hat{\mu}_{a}(t))+\sum_{i=1}^{M}l_{b_{i}}(\hat{\mu}_{b_{i}}(t)).$ (51)
Similarly, we have
$\log\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})$
$\qquad=\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}\bigg{[}l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg{]},$ (52)
where
$l_{x}(\mu^{\prime}_{x})=N_{x}(t)\bigg{(}\dot{A}^{-1}(\mu^{\prime}_{x})\hat{\mu}_{x}(t)-A(\dot{A}^{-1}(\mu^{\prime}_{x}))\bigg{)},\qquad\forall x\in\{a,b_{1},b_{2},\cdots,b_{M}\},$ (53)
is concave in $\mu^{\prime}_{x}$.
Now, we aim to find
$\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}\bigg{[}l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg{]}.$ (54)
The Lagrangian is
$L(\mu^{\prime}_{a},\mu^{\prime}_{b_{1}},\mu^{\prime}_{b_{2}},\cdots,\mu^{\prime}_{b_{M}},\beta_{1},\beta_{2},\cdots,\beta_{M})=l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})+\sum_{i=1}^{M}\beta_{i}(\mu^{\prime}_{a}-\mu^{\prime}_{b_{i}})$ (55)
where $\beta_{i}\geq 0$ for all $i\in[M]$.
Using the KKT conditions, we have
00 0 $=\frac{\partial L}{\partial\mu^{\prime}_{a}}=N_{a}(t)\big{(}\hat{\mu}_{a}(t)-\mu^{\prime}_{a}\big{)}\bigg{[}-\frac{\ddot{A}(\mu^{\prime}_{a})}{(\dot{A}(\mu^{\prime}_{a}))^{2}}\bigg{]}+\sum_{i=1}^{M}\beta_{i},$ (56)
00 0 $=\frac{\partial L}{\partial\mu^{\prime}_{b_{i}}}=N_{b_{i}}(t)\big{(}\hat{\mu}_{b_{i}}(t)-\mu^{\prime}_{b_{i}}\big{)}\bigg{[}-\frac{\ddot{A}(\mu^{\prime}_{b_{i}})}{(\dot{A}(\mu^{\prime}_{b_{i}}))^{2}}\bigg{]}-\beta_{i},\qquad\forall i\in[M],$ (57)
00 0 $=\beta_{i}(\mu^{\prime}_{a}-\mu^{\prime}_{b_{i}}),\qquad\forall i\in[M]$ (58)
$\mu^{\prime}_{a}$ $\geq\mu^{\prime}_{b_{i}},\qquad\forall i\in[M].$ (59)
Since $\beta_{i}\geq 0$ for all $i\in[M]$ from (56) and the fact that $\ddot{A}(\mu^{\prime}_{a})>0$ (exponential bandits), we have $\hat{\mu}_{a}(t)\geq\mu^{\prime}_{a}$. Now, from (58) if there exists $i\in[M]$ such that $\beta_{i}=0$, then from (57) we also have $\mu^{\prime}_{b_{i}}=\hat{\mu}_{b_{i}}(t)$. Hence, it follows from (59) we have
$\hat{\mu}_{a}(t)\geq\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{i}}=\hat{\mu}_{b_{i}}(t).$ (60)
Under the condition that $\hat{\mu}_{a}(t)\leq\hat{\mu}_{b_{i}}(t)),\forall i\in[M]$, it holds from (60) that
$\hat{\mu}_{a}(t)=\mu^{\prime}_{a}=\mu^{\prime}_{b_{i}}=\hat{\mu}_{b_{i}}(t).$ (61)
From (61) and (58), under the condition that $\hat{\mu}_{a}(t)\leq\hat{\mu}_{b_{i}}(t)),\forall i\in[M]$, we must have
$\mu^{\prime}_{a}=\mu^{\prime}_{b_{1}}=\mu^{\prime}_{b_{2}}=\cdots=\mu^{\prime}_{b_{M}}.$ (62)
Now, by summing all equations in (56) and (57) we obtain
$N_{a}(t)\big{(}\hat{\mu}_{a}(t)-\mu^{\prime}_{a}\big{)}\bigg{[}-\frac{\ddot{A}(\mu^{\prime}_{a})}{(\dot{A}(\mu^{\prime}_{a}))^{2}}\bigg{]}+\sum_{i=1}^{M}N_{b_{i}}(t)\big{(}\hat{\mu}_{b_{i}}(t)-\mu^{\prime}_{b_{i}}\big{)}\bigg{[}-\frac{\ddot{A}(\mu^{\prime}_{b_{i}})}{(\dot{A}(\mu^{\prime}_{b_{i}}))^{2}}\bigg{]}=0.$ (63)
From (62) and (63), under the condition that $\hat{\mu}_{a}(t)\geq\hat{\mu}_{b_{i}}(t)),\forall i\in[M]$, we obtain
$\mu^{\prime}_{a}=\mu^{\prime}_{b_{1}}=\mu^{\prime}_{b_{2}}=\cdots=\mu^{\prime}_{b_{M}}=\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t):=\frac{N_{a}(t)\hat{\mu}_{a}(t)+\sum_{i=1}^{M}N_{b_{i}}(t)\hat{\mu}_{b_{i}}(t)}{N_{a}(t)+\sum_{i=1}^{M}N_{b_{i}}(t)}.$ (64)
Hence, we have
$\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{2}}\wedge\cdots\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}\bigg{[}l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg{]}=l_{a}(\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))+\sum_{i=1}^{M}l_{b_{i}}(\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))$ (65)
By combining (62) and (65), under the condition $(\hat{\mu}_{a}(t)\leq\hat{\mu}_{b_{1}}(t))\wedge(\hat{\mu}_{a}(t)\leq\hat{\mu}_{b_{2}}(t))\wedge\cdots\wedge(\hat{\mu}_{a}(t)\leq\hat{\mu}_{b_{M}}(t))$, we obtain
$Z_{a;b_{1},b_{2},\cdots,b_{M}}(t)$ $=N_{a}(t)d(\hat{\mu}_{a}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))+\sum_{i=1}^{M}N_{b_{i}}(t)d(\hat{\mu}_{b_{i}}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)).$ (66)
We use the following stopping rule
$\tau_{\delta}$ $=\inf\bigg{\{}t\in\mathbb{N}:Z(t):=\max_{b_{1},b_{2},\cdots,b_{M}}\min_{a\in\mathcal{A}\setminus\{b_{1},b_{2},\cdots,b_{M}\}}Z_{a;b_{1},b_{2},\cdots,b_{M}}(t)>\beta(t,\delta)\bigg{\}},$ (67)
where $\beta(t,\delta)$ is an exploration rate to be adjusted.
The decoding rule is as following:
  • Let $(\hat{b}_{1},\hat{b}_{2},\cdots,\hat{b}_{M})=\text{argmax}_{b_{1},b_{2},\cdots,b_{M}}\min_{a\in\mathcal{A}\setminus\{b_{1},b_{2},\cdots,b_{M}\}}Z_{a;b_{1},b_{2},\cdots,b_{M}}(\tau_{\delta})$.
  • Choose $\hat{a}_{\tau}=\hat{b}_{i}$ with probability $1/M$ for all $i\in[M]$.
VI Choosing the threshold in the Stopping Rule
For simplicity, we only provide a proof for Bernoulli bandits (see (Garivier2016OptimalBA, )). The extension to other exponential bandits can be done by using the same methods as (Chambaz2005AMA, ).
Introducing, for $a,b_{1},b_{2},\cdots,b_{M}\in\mathcal{A}$,
$T_{a;b_{1},b_{2},\cdots,b_{M}}=\inf\{t\in\mathbb{N}:Z_{a;b_{1},b_{2},\cdots,b_{M}}(t)>\beta(t,\delta)\}.$
Then, we have
$\mathbb{P}_{\mu}\big{(}\tau_{\delta}<\infty,\hat{a}_{\tau_{\delta}}\notin\{1,2,\cdots,M\}\big{)}$
$\qquad\leq\mathbb{P}_{\mu}\Big{(}\exists\{b_{1},b_{2},\cdots,b_{M}\}\neq\{1,2,\cdots,M\},$
$\qquad\qquad\exists t\in\mathbb{N}:\min_{a\in\mathcal{A}\setminus\{b_{1},b_{2},\cdots,b_{M}\}}Z_{a;b_{1},b_{2},\cdots,b_{M}}(t)>\beta(t,\delta)\Big{)}.$ (68)
Now, for each tuple $(b_{1},b_{2},\cdots,b_{M})$ such that $\{b_{1},b_{2},\cdots,b_{M}\}\neq\{1,2,\cdots,M\}$, there exists
$a^{*}_{b_{1},b_{2},\cdots,b_{M}}\in\{1,2,\cdots,M\}\quad\text{such that}\quad a^{*}_{b_{1},b_{2},\cdots,b_{M}}\in\mathcal{A}\setminus\{b_{1},b_{2},\cdots,b_{M}\}.$
Then, we have
$\mu_{a^{*}_{b_{1},b_{2},\cdots,b_{M}}}\geq\max_{1\leq i\leq M}\mu_{b_{i}}.$
In addition, from (68) we obtain
$\mathbb{P}_{\mu}\big{(}\tau_{\delta}<\infty,\hat{a}_{\tau_{\delta}}\notin\{1,2,\cdots,M\}\big{)}$
$\qquad\leq\sum_{b_{1},b_{2},\cdots,b_{M}}\mathbb{P}_{\mu}\big{(}T_{a^{*}_{b_{1},b_{2},\cdots,b_{M}};b_{1},b_{2},\cdots,b_{M}}<\infty\big{)}.$ (69)
Now, it is sufficient to show that if
$\beta(t,\delta)=\log\left[K^{M}\left(\frac{4}{M+1}\right)^{(M+1)/2}t^{(M+1)/2}\frac{1}{\delta}\right]$
and
$\mu_{a}\geq\mu_{b_{1}}\wedge\mu_{a}\geq\mu_{b_{2}}\wedge\cdots\wedge\mu_{a}\geq\mu_{b_{M}},$
then
$\mathbb{P}_{\mu}\big{(}T_{a;b_{1},b_{2},\cdots,b_{M}}<\infty\big{)}\leq\frac{\delta}{K^{M}}.$
For such a tuple of arms, observe that on the event $\{T_{a;b_{1},b_{2},\cdots,b_{M}}=t\}$, time $t$ is the first moment when $Z_{a;b_{1},b_{2},\cdots,b_{M}}$ exceeds the threshold $\beta(t,\delta)$, which implies by definition that
$1\leq e^{-\beta(t,\delta)}\frac{\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{2}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}{\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{2}}\wedge\cdots\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}.$ (70)
It thus holds that
$\mathbb{P}_{\mu}\big{(}T_{a;b_{1},b_{2},\cdots,b_{M}}<\infty\big{)}$
$=\sum_{t=1}^{\infty}\mathbb{P}_{\mu}\big{(}T_{a;b_{1},b_{2},\cdots,b_{M}}=t\big{)}=\sum_{t=1}^{\infty}\mathbb{E}_{\mu}\left[\mathbbm{1}_{(T_{a;b_{1},b_{2},\cdots,b_{M}}=t)}\right]$ (71)
$\leq\sum_{t=1}^{\infty}e^{-\beta(t,\delta)}\mathbb{E}_{\mu}\Bigg{[}\mathbbm{1}_{(T_{a;b_{1},b_{2},\cdots,b_{M}}=t)}\frac{\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}{\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}\Bigg{]}$ (72)
$\leq\sum_{t=1}^{\infty}e^{-\beta(t,\delta)}\mathbb{E}_{\mu}\Bigg{[}\mathbbm{1}_{(T_{a;b_{1},b_{2},\cdots,b_{M}}=t)}\frac{\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}{p_{\mu_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}\Bigg{]}.$ (73)
Expanding the expectation into an integral over the sample space $\{0,1\}^{t}$, we get
$=\sum_{t=1}^{\infty}e^{-\beta(t,\delta)}\int_{\{0,1\}^{t}}\mathbbm{1}_{(T_{a;b_{1},b_{2},\cdots,b_{M}}=t)}(x_{1},x_{2},\cdots,x_{t})$
×$\quad\times\underbrace{\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})\prod_{i\in\mathcal{A}\setminus\{a,b_{1},\cdots,b_{M}\}}p_{\mu_{i}}(\underline{x}_{t}^{i})}_{\text{joint density under alternative parameters}}dx_{1}dx_{2}\cdots dx_{t}.$ (74)
Now, by using (Garivier2016OptimalBA, , Lemma 11), we have
$\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})$
$\qquad\leq 2\sqrt{n_{a}}\prod_{i=1}^{M}(2\sqrt{n_{b_{i}}})\,\text{kt}(\underline{x}_{t}^{a})\prod_{i=1}^{M}\text{kt}(\underline{x}_{t}^{b_{i}})$
$\qquad\leq 2^{M+1}\left(\frac{t}{M+1}\right)^{\frac{M+1}{2}}\text{kt}(\underline{x}_{t}^{a})\prod_{i=1}^{M}\text{kt}(\underline{x}_{t}^{b_{i}}),$ (75)
where $\text{kt}(\cdot)$ is the Krichevsky–Trofimov distribution (see (Krichevsky1981, )).
Therefore,
$\mathbb{P}_{\mu}\big{(}T_{a;b_{1},b_{2},\cdots,b_{M}}<\infty\big{)}$
$\leq\sum_{t=1}^{\infty}2^{M+1}\left(\frac{t}{M+1}\right)^{\frac{M+1}{2}}e^{-\beta(t,\delta)}\int_{\{0,1\}^{t}}\mathbbm{1}_{(T_{a;b_{1},b_{2},\cdots,b_{M}}=t)}(x_{1},\cdots,x_{t})$
×$\quad\times\text{kt}(\underline{x}_{t}^{a})\prod_{i=1}^{M}\text{kt}(\underline{x}_{t}^{b_{i}})\prod_{i\in\mathcal{A}\setminus\{a,b_{1},\cdots,b_{M}\}}p_{\mu_{i}}(\underline{x}_{t}^{i})\,dx_{1}\cdots dx_{t}.$ (76)
Defining $\tilde{\mathbb{P}}$ as the probability measure induced by the product of KT distributions for arms $a,b_{1},\cdots,b_{M}$ and the true distributions for all other arms, we have
$\mathbb{P}_{\mu}\big{(}T_{a;b_{1},b_{2},\cdots,b_{M}}<\infty\big{)}$ $\leq\sum_{t=1}^{\infty}2^{M+1}\left(\frac{t}{M+1}\right)^{\frac{M+1}{2}}e^{-\beta(t,\delta)}\tilde{\mathbb{P}}\big{(}T_{a;b_{1},b_{2},\cdots,b_{M}}=t\big{)}$
$\leq\frac{\delta}{K^{M}}\tilde{\mathbb{P}}\big{(}T_{a;b_{1},b_{2},\cdots,b_{M}}<\infty\big{)}\leq\frac{\delta}{K^{M}}.$ (77)
VII Sample Complexity Analysis
VII-A Almost-sure Upper Bound
Theorem 6.
Let $\mu$ denote a bandit model from an exponential family. Consider applying Chernoff’s stopping rule with confidence parameter
$\beta(t,\delta)=\log\left(\frac{Ct^{\alpha}}{\delta}\right),$
in conjunction with either the C-Tracking or D-Tracking sampling rule. Then, almost surely,
$\limsup_{\delta\to 0}\frac{\tau_{\delta}}{\log(1/\delta)}\leq T^{*}(\mu),$
where $\tau_{\delta}$ denotes the stopping time.
Proof.
For sufficiently large $t$, i.e., $t\geq t_{0}$, it holds that
$(\hat{\mu}_{1}(t)\geq\hat{\mu}_{a}(t))\vee(\hat{\mu}_{2}(t)\geq\hat{\mu}_{a}(t))\wedge\cdots\wedge(\hat{\mu}_{M}(t)\geq\hat{\mu}_{a}(t))$
for all $a\notin\{1,2,\ldots,M\}$.
Now observe that
$Z(t)$ $\geq\min_{a\notin\{1,\ldots,M\}}Z_{a;1,\ldots,M}(t)$ (78)
$=\min_{a\notin\{1,\ldots,M\}}\sum_{i=1}^{M}N_{i}(t)\,d\left(\hat{\mu}_{i}(t),\hat{\mu}_{a;1,\ldots,M}(t)\right)+N_{a}(t)\,d\left(\hat{\mu}_{a}(t),\hat{\mu}_{a;1,\ldots,M}(t)\right)$ (79)
$=t\cdot\min_{a\notin\{1,\ldots,M\}}\left(\sum_{i=1}^{M}\frac{N_{i}(t)}{t}+\frac{N_{a}(t)}{t}\right)$
$\quad\cdot I_{\frac{N_{1}(t)/t}{\sum_{i=1}^{M}N_{i}(t)/t+N_{a}(t)/t},\ldots,\frac{N_{M}(t)/t}{\sum_{i=1}^{M}N_{i}(t)/t+N_{a}(t)/t}}\left(\hat{\mu}_{1}(t),\ldots,\hat{\mu}_{M}(t),\hat{\mu}_{a}(t)\right).$ (80)
For $a\geq M+1$, the function
$(w,\lambda)\mapsto\left(\sum_{i=1}^{M}w_{i}+w_{a}\right)I_{\frac{w_{1}}{\sum_{i=1}^{M}w_{i}+w_{a}},\ldots,\frac{w_{M}}{\sum_{i=1}^{M}w_{i}+w_{a}}}(\lambda_{1},\ldots,\lambda_{M},\lambda_{a})$
is continuous at $(w^{*}(\mu),\mu)$. Therefore, for any $\varepsilon>0$, there exists $t_{1}\geq t_{0}$ such that for all $t\geq t_{1}$ and all $a\in\{M+1,\ldots,K\}$,
$\left(\sum_{i=1}^{M}\frac{N_{i}(t)}{t}+\frac{N_{a}(t)}{t}\right)$
$\quad\cdot I_{\frac{N_{1}(t)/t}{\sum_{i=1}^{M}N_{i}(t)/t+N_{a}(t)/t},\ldots,\frac{N_{M}(t)/t}{\sum_{i=1}^{M}N_{i}(t)/t+N_{a}(t)/t}}\left(\hat{\mu}_{1}(t),\ldots,\hat{\mu}_{M}(t),\hat{\mu}_{a}(t)\right)$
$\qquad\geq\frac{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}}{1+\varepsilon}\cdot I_{\frac{w_{1}^{*}}{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}},\ldots,\frac{w_{M}^{*}}{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}}}(\mu_{1},\ldots,\mu_{M},\mu_{a}).$ (81)
Hence, for $t\geq t_{1}$,
$Z(t)$ $\geq\frac{t}{1+\varepsilon}\cdot\min_{a\notin\{1,\ldots,M\}}\left(\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}\right)\cdot I_{\frac{w_{1}^{*}}{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}},\ldots,\frac{w_{M}^{*}}{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}}}(\mu_{1},\ldots,\mu_{M},\mu_{a})$
$=\frac{t}{(1+\varepsilon)T^{*}(\mu)}.$ (82)
Therefore,
$\tau_{\delta}$ $=\inf\{t\in\mathbb{N}:Z(t)\geq\beta(t,\delta)\}$ (83)
$\leq t_{1}\vee\inf\left\{t\in\mathbb{N}:\frac{t}{1+\varepsilon}\cdot\frac{1}{T^{*}(\mu)}\geq\log\left(\frac{Ct^{\alpha}}{\delta}\right)\right\},$ (84)
where
$C=K^{M}\left(\frac{4}{M+1}\right)^{(M+1)/2},\quad\alpha=\frac{M+1}{2}.$
Lemma 7.
Let $\alpha>0$ and $c_{1},c_{2}>0$ be such that
$A:=\frac{1}{\alpha}\log c_{2}+\log\left(\frac{\alpha}{c_{1}}\right)>\log e.$
Then there exists
$x\leq\frac{\alpha}{c_{1}}\left(A+\sqrt{2(A-1)}\right)$
such that
$c_{1}x\geq\log(c_{2}x^{\alpha}).$
In the case where $M$ is fixed (not a function of $\delta$), applying the lemma with $c_{1}=\frac{1}{(1+\varepsilon)T^{*}(\mu)}$ and $c_{2}=\frac{C}{\delta}$ yields
$A=\frac{1}{\alpha}\log\frac{C}{\delta}+\log\left(\alpha T^{*}(\mu)(1+\varepsilon)\right)>\log e$
as $\delta\to 0$. Therefore,
$\limsup_{\delta\to 0}\frac{\tau_{\delta}}{\log(1/\delta)}\leq(1+\varepsilon)T^{*}(\mu).$
Letting $\varepsilon\to 0$ concludes the proof.
VII-B Asymptotic Optimality in Expectation
Theorem 8.
Let $\mu$ be a bandit model from an exponential family. Using Chernoff’s stopping rule with
$\beta(t,\delta)=\log\left(\frac{Ct^{\alpha}}{\delta}\right),$
and the sampling rule C-Tracking or D-Tracking, we have
$\limsup_{\delta\to 0}\frac{\mathbb{E}_{\mu}[\tau_{\delta}]}{\log(1/\delta)}\leq T^{*}(\mu).$ (85)
Proof.
Following a similar strategy as in (Garivier2016OptimalBA, , Theorem 14), we simplify notation by assuming that the bandit means satisfy
$\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{M}>\mu_{M+1}\geq\mu_{M+2}\geq\cdots\geq\mu_{K}.$
Fix $\varepsilon>0$. By the continuity of $w^{*}(\mu)$, there exists $\xi=\xi(\varepsilon)\leq\frac{\mu_{M}-\mu_{M+1}}{4}$ such that the hypercube
$\mathcal{I}_{\varepsilon}:=[\mu_{1}-\xi,\mu_{1}+\xi]\times\cdots\times[\mu_{K}-\xi,\mu_{K}+\xi]$
satisfies, for all $\mu^{\prime}\in\mathcal{I}_{\varepsilon}$,
$\max_{a}\left|w_{a}^{*}(\mu^{\prime})-w_{a}^{*}(\mu)\right|\leq\varepsilon.$
In particular, if $\hat{\mu}(t)\in\mathcal{I}_{\varepsilon}$, then the top-$M$ empirical arms are exactly $\{1,2,\dots,M\}$.
Let $T\in\mathbb{N}$, define $h(T):=T^{1/4}$, and introduce the event
$\mathcal{E}_{T}:=\bigcap_{t=h(T)}^{T}\{\hat{\mu}(t)\in\mathcal{I}_{\varepsilon}\}.$
On the event $\mathcal{E}_{T}$, for $t\geq h(T)$ and any $a\notin\{1,2,\dots,M\}$, we have $\hat{\mu}_{i}(t)\geq\hat{\mu}_{a}(t)$ for all $i\in\{1,\dots,M\}$, and the Chernoff-type statistic satisfies
$\max_{b_{1},\dots,b_{M}}\min_{a\notin\{b_{1},\dots,b_{M}\}}Z_{a;b_{1},\dots,b_{M}}(t)$
$\qquad\geq\min_{a\notin\{1,\dots,M\}}Z_{a;1,\dots,M}(t)$ (86)
$\qquad=\min_{a\notin\{1,\dots,M\}}\left[N_{a}(t)d(\hat{\mu}_{a}(t),\hat{\mu}_{a;1,\dots,M}(t))+\sum_{i=1}^{M}N_{i}(t)d(\hat{\mu}_{i}(t),\hat{\mu}_{a;1,\dots,M}(t))\right]$ (87)
$\qquad=t\cdot g\left(\hat{\mu}(t),\left\{\frac{N_{i}(t)}{t}\right\}_{i=1}^{K}\right),$ (88)
where
$g(\mu^{\prime},w^{\prime}):=\min_{a\notin\{1,\dots,M\}}\left(\sum_{i=1}^{M}w^{\prime}_{i}+w^{\prime}_{a}\right)I_{\left\{\frac{w^{\prime}_{i}}{\sum_{j=1}^{M}w^{\prime}_{j}+w^{\prime}_{a}}\right\}_{i=1}^{M}}(\mu^{\prime}_{1},\dots,\mu^{\prime}_{M},\mu^{\prime}_{a}).$
Using (Garivier2016OptimalBA, , Lemma 20), for $T\geq T_{\varepsilon}$, define
$C^{*}_{\varepsilon}(\mu):=\inf_{\begin{subarray}{c}\|\mu^{\prime}-\mu\|_{\infty}\leq\xi(\varepsilon)\\ \|w^{\prime}-w^{*}(\mu)\|_{\infty}\leq 3(K-1)\varepsilon\end{subarray}}g(\mu^{\prime},w^{\prime}),$
so that on $\mathcal{E}_{T}$, for all $t\geq\sqrt{T}$,
$\max_{b_{1},\dots,b_{M}}\min_{a\notin\{b_{1},\dots,b_{M}\}}Z_{a;b_{1},\dots,b_{M}}(t)\geq tC^{*}_{\varepsilon}(\mu).$
Then, for $T\geq T_{\varepsilon}$, on $\mathcal{E}_{T}$,
$\min(\tau_{\delta},T)$ $\leq\sqrt{T}+\sum_{t=\sqrt{T}}^{T}\mathbbm{1}_{\tau_{\delta}>t}$ (89)
$\leq\sqrt{T}+\sum_{t=\sqrt{T}}^{T}\mathbbm{1}\left\{\max_{b_{1},\dots,b_{M}}\min_{a\notin\{b_{1},\dots,b_{M}\}}Z_{a;b_{1},\dots,b_{M}}(t)\leq\beta(t,\delta)\right\}$ (90)
$\leq\sqrt{T}+\sum_{t=\sqrt{T}}^{T}\mathbbm{1}\{tC^{*}_{\varepsilon}(\mu)\leq\beta(T,\delta)\}$ (91)
$\leq\sqrt{T}+\frac{\beta(T,\delta)}{C^{*}_{\varepsilon}(\mu)}.$ (92)
Define
$T_{0}(\delta):=\inf\left\{T\in\mathbb{N}:\sqrt{T}+\frac{\beta(T,\delta)}{C^{*}_{\varepsilon}(\mu)}\leq T\right\}.$
Then, for any $T\geq\max(T_{0}(\delta),T_{\varepsilon})$,
$\mathcal{E}_{T}\subset\{\tau_{\delta}\leq T\},$
and by (Garivier2016OptimalBA, , Lemma 19),
$\mathbb{P}_{\mu}(\tau_{\delta}>T)\leq\mathbb{P}(\mathcal{E}_{T}^{c})\leq BT\exp(-CT^{1/8}).$
Hence,
$\mathbb{E}_{\mu}[\tau_{\delta}]\leq T_{0}(\delta)+T_{\varepsilon}+\sum_{T=1}^{\infty}BT\exp(-CT^{1/8}).$
We now derive an upper bound on $T_{0}(\delta)$. For any $\eta>0$, define
$C(\eta):=\inf\left\{T\in\mathbb{N}:T-\sqrt{T}\geq\frac{T}{1+\eta}\right\},$
then
$T_{0}(\delta)$ $\leq C(\eta)+\inf\left\{T\in\mathbb{N}:\frac{1}{C^{*}_{\varepsilon}(\mu)}\log\left(\frac{CT^{\alpha}}{\delta}\right)\leq\frac{T}{1+\eta}\right\}$ (93)
$\leq C(\eta)+\inf\left\{T\in\mathbb{N}:\frac{C^{*}_{\varepsilon}(\mu)}{1+\eta}T\geq\log\left(\frac{CT^{\alpha}}{\delta}\right)\right\}.$ (94)
Applying Lemma 7 with $c_{1}=\frac{C^{*}_{\varepsilon}(\mu)}{1+\eta}$ and $c_{2}=\frac{C}{\delta}$ yields
$\lim_{\delta\to 0}\frac{\mathbb{E}_{\mu}[\tau_{\delta}]}{\log\frac{1}{\delta}}\leq\frac{1+\eta}{C^{*}_{\varepsilon}(\mu)}.$
Letting $\eta,\varepsilon\to 0$ and using continuity of $g$ and the definition of $w^{*}$ gives
$\lim_{\varepsilon\to 0}C^{*}_{\varepsilon}(\mu)=\frac{1}{T^{*}(\mu)},$
and thus,
$\lim_{\delta\to 0}\frac{\mathbb{E}_{\mu}[\tau_{\delta}]}{\log\frac{1}{\delta}}\leq T^{*}(\mu).$
VIII Conclusion
In this work, we revisited the problem of best-arm identification in stochastic multi-armed bandits under the fixed-confidence setting, with a particular focus on instances where multiple optimal arms may exist. While the Track-and-Stop algorithm has been widely regarded for its strong empirical performance, its theoretical guarantees in this more general setting have remained an open question.
Our primary contribution is a rigorous proof that Track-and-Stop achieves instance-optimal sample complexity even when the set of optimal arms is not a singleton. By carefully analyzing the generalized likelihood ratio statistics and aligning the stopping condition with information-theoretic lower bounds, we demonstrate that the algorithm retains its optimality across all problem instances—without requiring any modifications to its original formulation.
This result strengthens the theoretical foundation of likelihood ratio-based strategies for best-arm identification and affirms the robustness of Track-and-Stop beyond the classical unique-optimum setting. Looking ahead, an important direction for future work is to extend these insights to more structured bandit models, such as combinatorial or contextual bandits, and to develop adaptive algorithms that offer both optimal guarantees and improved empirical performance.
Appendix A Proof of Lemma 2
Let $P_{\mu}$ and $P_{\lambda}$ be distributions in a one-parameter exponential family. Then the KL divergence has the form:
$d(\mu,\lambda)=D(P_{\mu}\|P_{\lambda})=A(\theta(\lambda))-A(\theta(\mu))-(\theta(\lambda)-\theta(\mu))\mu.$ (95)
First, we show that for a fixed $\mu$, the function $d(\mu,\lambda)$ is convex in $\lambda$ for the common distributions above.
  • Gaussian Distribution (Known Variance):
    Let $X\sim\mathcal{N}(\mu,\sigma^{2})$. The natural parameter is $\theta(\lambda)=\lambda/\sigma^{2}$, and the log-partition function is $A(\theta)=\frac{\sigma^{2}\theta^{2}}{2}$.
    The KL divergence is:
    $d(\mu,\lambda)=\frac{(\mu-\lambda)^{2}}{2\sigma^{2}}.$
    This is a quadratic function in $\lambda$, and hence convex. In fact,
    $\frac{d^{2}}{d\lambda^{2}}d(\mu,\lambda)=\frac{1}{\sigma^{2}}>0.$
  • Bernoulli Distribution:
    Let $X\sim\text{Bernoulli}(\lambda)$ with $\lambda\in(0,1)$. The KL divergence is:
    $d(\mu,\lambda)=\mu\log\left(\frac{\mu}{\lambda}\right)+(1-\mu)\log\left(\frac{1-\mu}{1-\lambda}\right).$
    Taking the second derivative with respect to $\lambda$, we obtain:
    $\frac{d^{2}}{d\lambda^{2}}d(\mu,\lambda)=\frac{\mu}{\lambda^{2}}+\frac{1-\mu}{(1-\lambda)^{2}}.$
    Since $\lambda\in(0,1)$ and $\mu\in(0,1)$, this expression is strictly positive. Thus, $d(\mu,\lambda)$ is convex in $\lambda$.
  • Poisson Distribution:
    Let $X\sim\text{Poisson}(\lambda)$ with $\lambda>0$. The KL divergence is:
    $d(\mu,\lambda)=\lambda-\mu+\mu\log\left(\frac{\mu}{\lambda}\right).$
    Taking the first and second derivatives:
    $\frac{d}{d\lambda}d(\mu,\lambda)=1-\frac{\mu}{\lambda},\quad\frac{d^{2}}{d\lambda^{2}}d(\mu,\lambda)=\frac{\mu}{\lambda^{2}}>0.$
    Hence, $d(\mu,\lambda)$ is convex in $\lambda$ for Poisson distributions.
In all three cases—Gaussian, Bernoulli, and Poisson—the KL divergence $d(\mu,\lambda)$ is convex in the second argument $\lambda$, which is a useful property in many optimization and information-theoretic contexts.
Therefore, $\{d(\mu_{i},\lambda_{i})\}_{i=1}^{M}$ and $d(\mu_{a},\lambda_{a})$ are convex in their respective arguments.
Now, define
$f(\lambda_{1},\lambda_{2},\cdots,\lambda_{M},\lambda_{a})=\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a}).$ (96)
Since each term $d(\mu_{i},\lambda_{i})$ is convex in $\lambda_{i}$, and since $w_{1},w_{2},\cdots,w_{M},w_{a}>0$, their weighted sum is jointly convex in $(\lambda_{1},\lambda_{2},\cdots,\lambda_{M},\lambda_{a})$. This is a standard result from convex analysis.
Appendix B Proof of Lemma 7
First, we find $x$ such that
$c_{1}x=\log(c_{2}x^{\alpha})=\log c_{2}+\alpha\log x.$ (97)
This equation is equivalent to
$c_{1}x-\alpha\log x=\log c_{2}.$ (98)
Let $x=\frac{\alpha}{c_{1}}y$, then (98) is equivalent to
$y=\log y+A,$ (99)
or
$(-y)e^{-y}=-e^{-A}.$ (100)
Now, since $A\geq\log e$, we have $-e^{-A}\in[-\frac{1}{e},0)$. This means that
$y=-W_{-1}(-e^{-A})$ (101)
where $W_{-1}(\cdot)$ is the Lambert W function.
Now, by (Chatzigeorgiou2013BoundsOT, ), it holds that
$-1-\sqrt{2u}-u<W_{-1}(-e^{-u-1})<-1-\sqrt{2u}-\frac{2}{3}u,\qquad\forall u>0.$ (102)
It follows that
$y=-W_{-1}(-e^{-A})\leq A+\sqrt{2(A-1)},\qquad\text{if}\quad A>1.$ (103)
which leads to
$x\leq\frac{\alpha}{c_{1}}\big{(}A+\sqrt{2(A-1)}\big{)}.$ (104)
Acknowledgements
The authors sincerely thank Jonathan Scarlett for his valuable suggestions that helped improve the manuscript.